What is the probability that a lady has a red-haired gene? The probability of her mother having the gene is 2/3 and the probability of her father having the gene is 1/4.
Answer #1

The lady can have red haired gene if she gets the gene from either her mother of her father

She won't have the red haired gene if both her parents does not have the gene

P(mother having the gene) = 2/3

P(mother not having the gene) = 1 - 2/3 = 1/3

P(father having the gene) = 174

P(father not having the gene) = 1 - 1/4 = 3/4

P(a lady has a red-haired gene) = 1 - P(she does not have a red haired gene)

= 1 - 1/3 x 3/4

= 3/4